OpenStudy (anonymous):

Solve: x^2 + y^2 = 25 2x - y = 5

OpenStudy (anonymous):

from 2nd equation, y=2x-5

OpenStudy (anonymous):

substitute y=2x-5 into the 1st equation

OpenStudy (anonymous):

x^2+(4x^-20x+25)=25

OpenStudy (anonymous):

now re-arrange and you'll have 5x^2-20x=0

OpenStudy (anonymous):

so dividing by 5 gives x^2-4=0 x=2

OpenStudy (anonymous):

now put x=2 into y=2x-5 and you'll get y=-1